# Flatbands in tight-binding lattices with anisotropic potentials

### v1

@@ Abstract
We consider tight-binding models on Bravais lattices with anisotropic onsite potentials that vary along a given direction and are constant along the transverse one. Inspired by our previous work on flatbands in anti-$\mathcal{PT}$ symmetric Hamiltonians [1], we construct an anti-$\mathcal{PT}$ symmetric Hamiltonians with an $E=0$ flatband by tuning the hoppings and the shapes of potentials. This construction is illustrated for the square lattice with bounded and unbounded potentials. Unlike flatbands in short-ranged translationally invariant Hamiltonians, we conjecture that the considered $E=0$ flatbands do not host compact localized states. Instead the flatband eigenstates exhibit a localization transition along the potential direction upon increasing the potential strength for bounded potentials. For unbounded potentials flatband eigenstates are always localized irrespective of the potential strength.
@@ I Introduction
Flatbands are dispersionless bands in tight-binding Hamiltonians [2, 3], which are macroscopically degenerate, typically host compact localized states (CLS) and exhibit localization without disorder [4, 5]. Flatbands are attracting a lot of attention due to their extreme sensitivity to perturbations and emergent novel phases of matter with applications in various fields ranging from condensed matter systems to quantum technologies, e.g. high temperature superconductivity [6, 7, 8, 9, 10], subdimensional localization [11, 12], quantum chaos [13, 14], quantum hardware [15, 16, 17] etc.
@@
In translation invariant lattices, the construction of flatbands and their compact localized eigenstates have been studied extensively in the past years [18, 2, 19, 20, 21, 22, 23, 24, 25]. However flatbands can also exist in non-translationally invariance systems, like quasicrystals [26, 27, 28]. Another example is a tight-binding Hamiltonian on a $d$-dimensional lattice in the presence of a uniform Wannier-Stark field (linear potential). The field partially destroys translational invariance and all bands flat spectrum emerges [29]. These flatbands do not host CLS, but rather noncompact super-exponentially localized eigenstates. The Wannier-Stark flatbands were also studied in non-Bravais lattice settings [30, 1]. An interesting open issue is the existence of flatbands in other classes of non-translationally invariant Hamiltonians and localization properties of their eigenstates.
@@
In this work we propose Hamiltonians on Bravais lattices with an onsite potential that are antisymmetric under the joint action of reflection $\mathcal{P}$ and time reversal $\mathcal{T}$. The onsite potential varies along a given lattice direction and is constant in all the transverse directions. Such Hamiltonians have a single $E=0$ flatband unlike the DC field case [29], where all bands are flat. Our flatband construction is valid for all Bravais lattices and arbitrary range of hopping, as well as any orientation of the onsite potential that is a lattice vector.
@@
We provide several examples for the tight-binding Hamiltonians on the square lattice with bounded potentials: quasi-periodic Aubry-André-like potential [31] and an inverse trigonometric potential, and an unbounded potential. In contrast to the conventional flatbands which feature compact localized states and do not support delocalized eigenstates, here by analyzing Bloch Hamiltonian and inverse participation ratio for flatband eigenstates we demonstrate a localization-delocalization transition along the field can happen. This transition is consistent with the observations in Refs. 32, 33. For the example of an unbounded polynomial potential we find that the eigenstates of the flatband are always localized.
@@
The article is organized as follows: In Sec. II we define the generic Hamiltonian with anisotropic onsite potential. In Sec. III we define anti-symmetry operator and analyze possible constraint on a tight-binding Hamiltonian with arbitrary long-range hopping in simple square lattices—which supports $E=0$ flatband—we provide the abstract existential proof of the flatband in a general $d$-dimensional Bravais lattice in Appendix A. In Sec. IV we provide numerical evidences of the flatband and possible localization-delocalization transition with both bounded and unbounded potentials. We concluded in Sec. V.
@@ II Model with an anisotropic potential
We consider a square lattice with sites labelled by an integer vector $\vec{n}=(n_{1},n_{2})\in\mathbb{Z}^{2}$ and a tight-binding Hamiltonian with an onsite potential:
@@
We assume that the potential $V(n_{1},n_{2})$ varies along a lattice vector $\vec{z}$ and is constant in the transverse direction $\vec{w}$, which is also a lattice vector for the square lattice. The present setting can be considered as a generalization of the DC field case, where the DC field was applied along the direction $\vec{z}$, that we considered previously in Ref. 29. Since both $\vec{z},\vec{w}$ are lattice vectors we can express them as follows
$z=\alpha_{1}n_{1}+\alpha_{2}n_{2},\quad w=\alpha_{2}n_{1}-\alpha_{1}n_{2}$ (2)
where $|\alpha_{1}|$ and $|\alpha_{2}|$ are either mutually prime numbers or tuples $\{(0,\pm 1),(\pm 1,0)\}$ 111 $(\alpha_{1},\alpha_{2})$ are related to the DC field direction for Wannier-Stark system [29]. In what follows we use either the tuple $(\alpha_{1},\alpha_{2})$ or the decomposition of $z$ to denote the direction in which the potential varies in space. Thanks to the Bézout’s identity both $z$ and $w$ are ensured to take integer values only. However there are constraints on the allowed values of $w$ depending on the value of $z$ and vice versa, except for the simple cases when one of $\alpha_{1,2}$ is zero [29]. For example if the potential changes only along the main diagonal of the square lattice, $(1,1)$, as shown in Fig. 1 then $z=n_{1}+n_{2},w=n_{1}-n_{2}=z-2n_{2}$ and $w$ can only take even (odd) values for even (odd) $z$. Therefore it is more convenient to work with $(z,\eta)$, $\eta\equiv n_{2}$ rather than $(z,w)$ in this case. The proper way of defining $\eta$ was defined in Ref. 29 for a generic 2D Bravais lattice and generic tuples $(\alpha_{1},\alpha_{2})$.
Figure 1: Blue spheres indicate lattice sites, black solid lines are hopping connections, red arrowed lines are for the direction $z$ and lattice basis vectors $n_{1},n_{2}$. Square hopping network with tilted field: the field is along the diagonal with respect to the unit cell basis vectors
The translation invariance of the Hamiltonian along $w$ implies the invariance in $\eta$. Therefore the Hamiltonian (1) can be block diagonalized using the Fourier transform with an associated momentum $k$:
@@
where $l_{z}=\alpha_{1}l_{1}+\alpha_{2}l_{2}$ is the hopping step in the $z$ direction, and $l_{\eta}$ is the step of the $z$-independent coordinate $\eta$ along the hopping direction. The value of $l_{\eta}$ is a linear function of $\vec{l}$, its exact form depends on the details of the Hamiltonian [29] and is provided below for several examples. We have also assumed either periodic or infinite system along $\vec{w}$ to properly define the momentum $k$.
@@
III Anti-symmetry induced flatband
We define a time reversal operator $\mathcal{T}:\ket{\psi}\mapsto\ket{\psi^{*}}$ (with respect to the position basis) and a parity operator $\mathcal{P}$:
$\mathcal{P}=\sum_{\vec{n}}(-1)^{z}\ket{-\vec{n}}\bra{\vec{n}}$ | (4)
which reflects the lattice site vector $\vec{n}$ with respect to the origin $z=0$ and adds a $z$-dependent phase. We require an anti-symmetry of the full Hamiltonian (1) under the joint action of $\mathcal{P}$ and $\mathcal{T}$: $\mathcal{R}=\mathcal{T}\cdot\mathcal{P}$
@@ III Anti-symmetry induced flatband
This requires the potential $V$ to be odd with respect to the reflection of $\vec{z}$ (4):
@@
and imposes a non-trivial constraint on the hoppings
@@
The Hermiticity condition, $[t(-\vec{l})]^{*}=t(\vec{l})$, implies that only odd $l_{z}$ are allowed. As a consequence $l_{z}\neq 0$ and no equipotential hopping is allowed—similarly to the Wannier-Stark flatband case [29]. The absence of equipotential hopping holds for any Bravais lattice model—see Eq. (26) in Appendix A. There might exist other possible choices of the $R$.
Now we observe that the momentum |$\ket{k}\mapsto\ket{-k}$ under the action of either $\mathcal{P}$ or $\mathcal{T}$, but it is invariant under the joint action of $\mathcal{T}\cdot\mathcal{P}$. Therefore the Hamiltonian $\mathcal{H}$ is anti-symmetric under the action of $\mathcal{T}\cdot\mathcal{P}$ for each momentum $k$ independently
@@
Therefore the eigenvalues of $\mathcal{H}$ come in pairs $(E(k),-E(k))$ with corresponding eigenvectors (|$(\ket{\psi(k)}$, $\matrixelement{k}{\mathcal{T}\cdot\mathcal{P}}{k}\ket{\psi(k)})$ for each momentum $k$. For an odd number of available $z$ values, the Hamiltonian $\mathcal{H}(k)$ has an odd number of eigenvalues for each $k$. Consequently there is necessarily zero eigenvalues, $E(k)=0$, for all momenta $k$, e.g. a flatband, which is located at the middle of the spectrum. This is similar to the case of chiral flatbands [30] and the requirement of the odd number of sublattices for the existence of anti-$\mathcal{PT}$ symmetric flatband for non-Bravais lattices [1]. For a finite system of size $L$ along the $z$ direction, this implies an odd number of equipotential lines (value of $z$ is fixed on each of the lines): $z\in[-(L-1)/2,(L-1)/2]$. This implies an odd number of energy bands $E(k)$.
Two examples of the potential $V$
We provide two example settings on the square lattice, that satisfy the above requirements.
One example is shown in Fig. 1: the potential only varies along the main diagonal of the square lattice, $z=n_{1}+n_{2}$. For the nearest neighbor hopping we have $(l_{1},l_{2})\in\{(0,\pm 1),(\pm 1,0)\}$. Therefore $l_{z}\in\{1,-1\}$ and there is no equipotential hopping [Eq. (7)], i.e. no hopping is present along the anti-diagonal, $w=n_{1}-n_{2}$ in the absence of hopping along $z$. The corresponding Hamiltonian reads
@@
In this case the anti-diagonal coordinate $w$ is a function of $z$ and contains a $z$-independent part. We define $\eta:=n_{2}$ as the independent coordinate [29]:
@@
The Hamiltonian $\mathcal{H}_{1,1}$ reads in terms of the $(z,\eta)$ coordinates
@@
Note that for $kl=\pm\pi$ the hopping part vanishes leading to compact localized structure of the eigenstates for all eigenvalues.
Figure 2: Blue spheres indicate the square lattice sites. The indices are $(n_{1},n_{2})$ with the corresponding basis vector directions indicated by red arrow. The red arrow with the label $z$ indicates the $z$ direction.
@@
In the momentum space (with respect to $\eta$) the Hamiltonian reads
@@ IV Numerical results
We now verify the above construction, Eqs. (12) and (III), for several example potentials. We consider quasi-periodic potential which is easily realizable in the state-of-art experimental devices, and a trigonometric and a polynomial potentials [35]. The Hamiltonians are defined on the square lattice wrapped around a cylinder in the $(z,k)$-space assuming periodicity along the $k$-direction. Along the $z$ direction (the potential) we assume the cylinder large but finite. We denote the size along $z$ as $L$ and refer to it as the system size in what follows. The momentum-space Hamiltonians given by Eqs. (12) and (III) are diagonalized numerically.
The macroscopic degeneracy of the flatband eigenstates |$\ket{\psi_{E=0}(k)}$ is a challenge: any linear combination of eigenstates is also an eigenstate. Translationally invariant Hamiltonians with a flatband host CLS [36], which provide a convenient basis for analysis [25]. However we conjecture, that flatbands under consideration do not have CLS, as could potentially be shown by extending the same argument used for the Wannier-Stark flatbands [29]. We choose the $z,k$ basis, that is natural for the model, to study the flatband eigenstates and their dependence on the potential strength $\lambda$. This implies in particular that eigenstates in the direction transverse to $z$ are always delocalized for this “choice of flatband eigenstates”. To quantify localization of the states along $z$ we compute the participation entropy
@@ IV.1 Aubry-André potential
Figure 3: (a)-(b) Band structure containing flatband at $E=0$ for the Hamiltonian in Eq. (11) with $\lambda=5$. (c)-(d) Band structure containing flatband at $E=0$ for the Hamiltonian in Eq. (14) with $\lambda=1.6$. Subfigures (b) and (d) are the spectra shown in subfigures (a) and (c) respectively, magnified around the $E=0$ flatband. $201$ lattice sites along $z$ are considered with open boundary condition. The reference (center of reflection for parity operation) site $\vec{n}=\vec{0}$ is set at the center of the lattice. $\beta=\left(\sqrt{5}+1\right)/2$.
We consider the Aubry-André (AA) potential of strength $\lambda$ and irrational parameter $\beta$
@@
for the two Hamiltonians in Eqs. (12) and (III) varying along directions: $(1,1):z=n_{1}+n_{2}$ and $(3,1):z=3n_{1}+n_{2}$ respectively. Using exact diagonalization we computed the spectrum for potential strength $\lambda=5$ and $\beta=(1+\sqrt{5})/2$, and system size $L=201$ with open boundary condition along the $z$ direction as shown in Fig. 3. The spectrum is symmetric with respect to $E=0$ and a flatband is present at $E=0$ due to the choice of the odd system size $L$.
Figure 4: $\mathcal{S}(k,\lambda)/\ln(L)$ as a function of $k,\lambda$ for the $E=0$ flatband eigenstates for quasi-periodic potential (17). System size $L=201$ along the field. Subfigures (a) and (b) correspond to the potential directions as in Figs. 1 and 2, respectively.
Figure 5: Participation entropy in units of $\ln(L)$ as a function of the potential strength $\lambda$ (17), for different system sizes $L$ along the z-direction. Potential directions (a): $(1,1),z=n_{1}+n_{2}$ and (b): $(3,1),z=3n_{1}+n_{2}$. Solid lines correspond to momentum $k=0$ while dashed lines correspond to $k=3\pi/4$. Note that for the direction (3, 1) the curves for the two momenta are barely distinguishable.
The Hamiltonian $\mathcal{H}$ is diagonalized for a discrete set of $k$ and several values of $\lambda$, the participation entropy (16) is then computed from the set of eigenvalues and eigenvectors $\{E(k),\ket{\psi(k)}\}$. The violet (dark) regions in Fig. 4 correspond to the flatband eigenstates localized along the field while yellow/green (bright) regions correspond to eigenstates delocalized along the $z$ direction. We observe a sharp, $k$-dependent transition from delocalized (yellow/green) to localized (violet) eigenstates with increasing potential strength $\lambda$. Similar results are obtained for other system sizes $L$, as shown in Fig. 5. The transition for the direction $(1,1)$ i.e., $z=n_{1}+n_{2}$ matches well with the analytical prediction [32], up to additional phase factors (see Appendix B).
@@
Figure 6: Profiles of the flatband eigenstates as functions of $z$ for different momenta and different orientation of the potential at fixed $\lambda=10$ (17), where all $E=0$ are localized. The direction $(\alpha_{1},\alpha_{2})=(3,1)$ leads to larger localization volume than that for the direction $(\alpha_{1},\alpha_{2})=(1,1)$ for the same set of parameters.
We have considered a finite irrational parameter $\beta=(\sqrt{5}+1)/2$. Different behavior, e.g. pronounced intermediate localization completely hiding exponential decay, might emerge upon decreasing the value of $\beta$, extending the results of Ref. 37.

### v2

@@ Abstract
We consider tight-binding models on Bravais lattices with anisotropic onsite potentials that vary along a given direction and are constant along the transverse one. Inspired by our previous work on flat bands in anti-$\mathcal{PT}$ symmetric Hamiltonians [Mallick et al., Phys. Rev. A 105, L021305 (2022)], we construct an anti-$\mathcal{PT}$ symmetric Hamiltonians with an $E=0$ flat band by tuning the hoppings and the shapes of potentials. This construction is illustrated for the square lattice with bounded and unbounded potentials. Unlike flat bands in short-ranged translationally invariant Hamiltonians, we conjecture that the considered $E=0$ flat bands do not host compact localized states. Instead the flat-band eigenstates exhibit a localization transition along the potential direction upon increasing the potential strength for bounded potentials. For unbounded potentials flat-band eigenstates are always localized irrespective of the potential strength.
@@ I Introduction
Flat bands are dispersionless bands in tight-binding Hamiltonians [1, 2], which are macroscopically degenerate, typically host compact localized states (CLSs), and exhibit localization without disorder [3, 4]. Flat bands attract a lot of attention due to their extreme sensitivity to perturbations and emergent novel phases of matter with applications in various fields ranging from condensed matter systems to quantum technologies, e.g., high temperature superconductivity [5, 6, 7, 8, 9], subdimensional localization [10, 11], quantum chaos [12, 13], quantum hardware [14, 15, 16], etc.
@@
In translation invariant lattices, the construction of flat bands and their compact localized eigenstates has been studied extensively in the past years [17, 1, 18, 19, 20, 21, 22, 23, 24]. However flat bands can also exist in nontranslationally invariance systems like quasicrystals [25, 26, 27]. Another example is a tight-binding Hamiltonian on a $d$-dimensional Bravais lattice in the presence of a uniform Wannier-Stark field (linear potential). The field partially destroys translational invariance and all-bands-flat spectrum emerges [28]. These flat bands do not host CLSs, but rather noncompact superexponentially localized eigenstates. The Wannier-Stark flat bands were also studied in non-Bravais lattice settings [29, 30]. An interesting open issue is the existence of flat bands in other classes of nontranslationally invariant Hamiltonians and localization properties of their eigenstates.
@@
In this paper, we propose Hamiltonians on Bravais lattices with an on-site potential that are antisymmetric under the joint action of reflection $\mathcal{P}$ and time reversal $\mathcal{T}$. The on-site potential varies along a given lattice direction and is constant in all the transverse directions. Such Hamiltonians have a single $E=0$ flat band unlike the dc field case [28], where all bands are flat. Our flat-band construction is valid for all Bravais lattices and arbitrary range of hopping as well as any on-site potential, so long as it varies along a direction that is a lattice vector.
@@
We provide several examples for the tight-binding Hamiltonians on the square lattice with bounded potentials: quasi-periodic Aubry-André (AA)-like potential [31], an inverse trigonometric potential, and an unbounded potential. In contrast to the conventional flat bands which feature compact localized states and do not support delocalized eigenstates, here, by analyzing Bloch Hamiltonian and inverse participation ratio for flat-band eigenstates, we demonstrate that a localization-delocalization transition along the field can happen. This transition is consistent with the observations in Refs. [32, 33]. For the example of an unbounded polynomial potential, we find that the eigenstates of the flat band are always localized.
@@
The article is organized as follows: In Sec. II, we define the generic Hamiltonian with anisotropic on-site potential. In Sec. III, we define the antisymmetry operator and analyze possible constraint on a tight-binding Hamiltonian with arbitrary long-range hopping in simple square lattices—which supports an $E=0$ flat band—and we provide the abstract existential proof of the flat band in a general $d$-dimensional Bravais lattice in Appendix A. In Sec. IV, we provide numerical evidences of the flat band and possible localization-delocalization transition with both bounded and unbounded potentials. We concluded in Sec. V.
@@ II Model with an anisotropic potential
We consider a square lattice with sites labeled by an integer vector $\vec{n}=(n_{1},n_{2})\in\mathbb{Z}^{2}$ and a tight-binding Hamiltonian with an on-site potential:
@@
We assume that the potential $V(n_{1},n_{2})$ varies along a lattice vector $\vec{z}$ and is constant in the transverse direction $\vec{w}$, which is also a lattice vector for the square lattice. The present setting can be considered as a generalization of the dc field case, where the dc field was applied along the direction $\vec{z}$, that we considered previously in Ref. [28]. Since both $\vec{z},\vec{w}$ are lattice vectors, we can express them as follows:
$z=\alpha_{1}n_{1}+\alpha_{2}n_{2},\quad w=\alpha_{2}n_{1}-\alpha_{1}n_{2},$ (2)
where $|\alpha_{1}|$ and $|\alpha_{2}|$ are either mutually prime numbers or tuples $\{(0,\pm 1),(\pm 1,0)\}$ 111$(\alpha_{1},\alpha_{2})$ are related to the DC field direction for Wannier-Stark system [28]. In what follows, we use either the tuple $(\alpha_{1},\alpha_{2})$ or the decomposition of $z$ to denote the direction in which the potential varies in space. Thanks to the Bézout’s identity, both $z$ and $w$ are ensured to take integer values only. However, there are constraints on the allowed values of $w$ depending on the value of $z$ and vice versa, except for the simple cases when one of $\alpha_{1,2}$ is zero [28]. For example, if the potential changes only along the main diagonal of the square lattice, $(1,1)$, as shown in Fig. 1, then $z=n_{1}+n_{2},w=n_{1}-n_{2}=z-2n_{2}$ and $w$ can only take even (odd) values for even (odd) $z$. Therefore, it is more convenient to work with $(z,\eta)$, $\eta\equiv n_{2}$ rather than $(z,w)$ in this case. The proper way of defining $\eta$ was defined in Ref. [28] for a generic 2D Bravais lattice and generic tuples $(\alpha_{1},\alpha_{2})$.
Figure 1: Blue spheres indicate lattice sites, black solid lines are hopping connections, red arrowed lines are for the direction $z$ and lattice basis vectors $n_{1},n_{2}$. Square hopping network with tilted field: the field is along the diagonal with respect to the unit cell basis vectors.
The translation invariance of the Hamiltonian along $w$ implies the invariance in $\eta$. Therefore, the Hamiltonian (1) can be block diagonalized using the Fourier transform with an associated momentum $k$:
@@
where $l_{z}=\alpha_{1}l_{1}+\alpha_{2}l_{2}$ is the hopping step in the $z$ direction, and $l_{\eta}$ is the step of the $z$-independent coordinate $\eta$ along the hopping direction. The value of $l_{\eta}$ is a linear function of $\vec{l}$; its exact form depends on the details of the Hamiltonian [28] and is provided below for several examples. We have also assumed either periodic or infinite system along $\vec{w}$ to properly define the momentum $k$.
@@
III Antisymmetry-induced flat band
We define a time-reversal operator $\mathcal{T}:\ket{\psi}\mapsto\ket{\psi^{*}}$ (with respect to the position basis) and a parity operator $\mathcal{P}$,
$\mathcal{P}=\sum_{\vec{n}}(-1)^{z}\ket{-\vec{n}}\bra{\vec{n}},$ (4)
which reflects the lattice site vector $\vec{n}$ with respect to the origin $\vec{n}=\vec{0}$ and adds a $z$-dependent phase. We require an antisymmetry of the full Hamiltonian (1) under the joint action of $\mathcal{P}$ and $\mathcal{T}$: $\mathcal{R}=\mathcal{T}\cdot\mathcal{P}$:
@@ III Antisymmetry-induced flat band
This requires the potential $V$ to be odd with respect to the reflection of $\vec{z}$ (4),
@@
and imposes a nontrivial constraint on the hoppings:
@@
The Hermiticity condition, $[t(-\vec{l})]^{*}=t(\vec{l})$, implies that only odd $l_{z}$ are allowed. As a consequence, $l_{z}\neq 0$ and no equipotential hopping is allowed—similarly to the Wannier-Stark flat-band case [28]. The absence of equipotential hopping holds for any Bravais lattice model—see Eq. (26) in Appendix A. There might exist other possible choices of the $R$.
Now we observe that the momentum |$\ket{k}\mapsto\ket{-k}$ under the action of either $\mathcal{P}$ or $\mathcal{T}$, but it is invariant under the joint action $\mathcal{T}\cdot\mathcal{P}$. Therefore, the Hamiltonian $\mathcal{H}$ is anti-symmetric under the action of $\mathcal{T}\cdot\mathcal{P}$ for each momentum $k$ independently:
@@
Therefore, the eigenvalues of $\mathcal{H}$ come in pairs $(E(k),-E(k))$ with corresponding eigenvectors (|$(\ket{\psi(k)}$, $\matrixelement{k}{\mathcal{T}\cdot\mathcal{P}}{k}\ket{\psi(k)})$ for each momentum $k$. For an odd number of available $z$ values, the Hamiltonian $\mathcal{H}(k)$ has an odd number of eigenvalues for each $k$. Consequently, there is necessarily zero eigenvalues, $E(k)=0$, for all momenta $k$, i.e., a flat band, which is located at the middle of the spectrum. This is similar to the case of chiral flat bands [29] and the requirement of the odd number of sublattices for the existence of anti-$\mathcal{PT}$ symmetric flat bands for non-Bravais lattices [30]. For a finite system of size $L$ along the $z$ direction, this implies an odd number of equipotential lines (value of $z$ is fixed on each of the lines): $z\in[-(L-1)/2,(L-1)/2]$. This implies an odd number of energy bands $E(k)$.
III.1 Two examples of the potential $V$
We provide two example settings on the square lattice that satisfy the above requirements.
One example is shown in Fig. 1: the potential only varies along the main diagonal of the square lattice, $z=n_{1}+n_{2}$. For the nearest-neighbor hopping, we have $(l_{1},l_{2})\in\{(0,\pm 1),(\pm 1,0)\}$. Therefore, $l_{z}\in\{1,-1\}$ and there is no equipotential hopping [Eq. (7)], i.e., no hopping is present along the anti-diagonal, $w=n_{1}-n_{2}$ in the absence of hopping along $z$. The corresponding Hamiltonian reads
@@ III.1 Two examples of the potential $V$
In this case, the antidiagonal coordinate $w$ is a function of $z$ and contains a $z$-independent part. We define $\eta:=n_{2}$ as the independent coordinate [28]:
@@
The Hamiltonian $\mathcal{H}_{1,1}$ reads in terms of the $(z,\eta)$ coordinates,
@@
Note that for $kl=\pm\pi$, the hopping part vanishes, leading to compact localized structure of the eigenstates for all eigenvalues.
Figure 2: Blue spheres indicate the square lattice sites. The indices are $(n_{1},n_{2})$ with the corresponding basis vector directions indicated by the red arrow. The red arrow with the label $z$ indicates the $z$ direction.
@@
In the momentum space (with respect to $\eta$), the Hamiltonian reads
@@ IV Numerical results
We now verify the above construction, Eqs. (12) and (III.1), for several example potentials. We consider quasi-periodic potential, which is easily realizable in state-of-the-art experimental devices, and trigonometric and polynomial potentials [35]. The Hamiltonians are defined on the square lattice wrapped around a cylinder in the $(z,k)$ space, assuming periodicity along the $k$ direction. Along the $z$ direction (the potential), we assume the cylinder large but finite. We denote the size along $z$ as $L$ and refer to it as the system size in what follows. The momentum-space Hamiltonians given by Eqs. (12) and (III.1) are diagonalized numerically.
The macroscopic degeneracy of the flat-band eigenstates |$\ket{\psi_{E=0}(k)}$ is a challenge: any linear combination of eigenstates is also an eigenstate. Translationally invariant Hamiltonians with a flat band host CLS [36], which provide a convenient basis for analysis [24]. However, we conjecture that flat bands under consideration do not have CLS, as could potentially be shown by extending the same argument used for the Wannier-Stark flat bands [28]. We choose the $z,k$ basis, that is natural for the model, to study the flat-band eigenstates and their dependence on the potential strength $\lambda$. This implies, in particular, that eigenstates in the direction transverse to $z$ are always delocalized for this choice of flatband eigenstates. To quantify localization of the states along $z$, we compute the participation entropy:
@@ IV.1 Aubry-André potential
Figure 3: (a)-(b) Band structure containing flat band at $E=0$ for the Hamiltonian in Eq. (11) with $\lambda=5$. (c)-(d) Band structure containing flat band at $E=0$ for the Hamiltonian in Eq. (14) with $\lambda=1.6$. (b), (d) The spectra shown in subfigures (a) and (c), respectively, magnified around the $E=0$ flat band. $201$ lattice sites along $z$ are considered with open boundary condition. The reference (center of reflection for parity operation) site $\vec{n}=\vec{0}$ is set at the center of the lattice. $\beta=\left(\sqrt{5}+1\right)/2$.
We consider the AA potential of strength $\lambda$ and irrational parameter $\beta$,
@@
for the two Hamiltonians in Eqs. (12) and (III.1) varying along directions $(1,1):z=n_{1}+n_{2}$ and $(3,1):z=3n_{1}+n_{2}$, respectively. Using exact diagonalization, we computed the spectrum for potential strengths $\lambda=5$ and $\lambda=1.6$ for directions (1,1) and (3,1), respectively, $\beta=(1+\sqrt{5})/2$, and system size $L=201$ with open boundary condition along the $z$ direction as shown in Fig. 3. The spectrum is symmetric with respect to $E=0$ and a flat band is present at $E=0$ due to the choice of the odd system size $L$.
Figure 4: $\mathcal{S}(k,\lambda)/\ln(L)$ as a function of $k,\lambda$ for the $E=0$ flat-band eigenstates for quasi-periodic potential (17). System size $L=201$ along the field. (a) and (b) correspond to the potential directions as in Figs. 1 and 2, respectively.
Figure 5: Participation entropy in units of $\ln(L)$ as a function of the potential strength $\lambda$ (17) for different system sizes $L$ along the z-direction. Potential directions (a) $(1,1),z=n_{1}+n_{2}$ and (b) $(3,1),z=3n_{1}+n_{2}$. Solid lines correspond to momentum $k=0$ while dashed lines correspond to $k=3\pi/4$. Note that for the direction (3, 1), the curves for the two momenta are barely distinguishable.
The Hamiltonian $\mathcal{H}$ is diagonalized for a discrete set of $k$ and several values of $\lambda$, the participation entropy (16) is then computed from the set of eigenvalues and eigenvectors $\{E(k),\ket{\psi(k)}\}$. The violet (dark) regions in Fig. 4 correspond to the flat-band eigenstates localized along the field while yellow/green (bright) regions correspond to eigenstates delocalized along the $z$ direction. We observe a sharp, $k$-dependent transition from delocalized (yellow/green) to localized (violet) eigenstates with increasing potential strength $\lambda$. Similar results are obtained for other system sizes $L$, as shown in Fig. 5. The transition for the direction $(1,1)$, i.e., $z=n_{1}+n_{2}$ matches well with the analytical prediction [32], up to additional phase factors (see Appendix B).
@@
Figure 6: Profiles of the flat-band eigenstates as functions of $z$ for different momenta and different orientation of the potential at fixed $\lambda=10$ (17), where all $E=0$ are localized along $z$. The direction $(\alpha_{1},\alpha_{2})=(3,1)$ leads to larger localization volume than that for the direction $(\alpha_{1},\alpha_{2})=(1,1)$ for the same set of parameters.
We have considered a finite irrational parameter $\beta=(\sqrt{5}+1)/2$. Different behaviors, e.g., pronounced intermediate localization completely hiding exponential decay, might emerge upon decreasing the value of $\beta$, extending the results of Ref. [37].